# Consequences of a Two-Time Relativistic Bohmian Model

### v16

@@ II Relativistic generalization issue for the quantum potential in case of constant momentum
$\frac{\partial S}{\partial t}+\frac{(\vec{\nabla}S){}^{2}}{2m}\,+V-\frac{\hbar{}^{2}}{2m}\,\frac{\nabla{}^{2}R}{R}=0$ (2)
$\frac{\partial R^{2}}{\partial t}+\vec{\nabla}\cdot(R^{2}\frac{\vec{\nabla}S}{m})=0$ (3)
@@
$H=\frac{p{}^{2}}{2m}+V-\frac{\hbar{}^{2}}{2m}\,\frac{\nabla{}^{2}R}{R}$ (4)
@@
$V_{Q}\equiv-\frac{\hbar{}^{2}}{2m}\,\frac{\nabla{}^{2}R}{R}$ (5)
@@
$\frac{d\vec{p}}{dt}=-\vec{\nabla}V-\vec{\nabla}V_{Q}$ (6)
@@
$\frac{\nabla{}^{2}\psi}{\psi}=\frac{\nabla{}^{2}R}{R}+2\,\frac{i}{\hbar}\,\frac{\vec{\nabla}R}{R}\,\cdot\,\vec{p}-\frac{p^{2}}{\hbar{}^{2}}$ (7)
@@
$\nabla{}^{2}\psi-\frac{1}{c^{2}}\,\frac{\partial^{2}\psi}{\partial t^{2}}=\frac{m{}^{2}c^{2}}{\hbar{}^{2}}\,\psi$ (8)
@@
$\frac{\nabla{}^{2}R}{R}+2\,\frac{i}{\hbar}\,\frac{\vec{\nabla}R}{R}\,\cdot\,\vec{p}=\frac{m{}^{2}\gamma^{2}c^{2}}{\hbar{}^{2}}+\frac{1}{c^{2}\psi}\,\frac{\partial^{2}\psi}{\partial t^{2}}$ (9)
@@
$\frac{1}{\psi}\frac{\partial^{2}\psi}{\partial t^{2}}=\frac{1}{R}\frac{\partial^{2}R}{\partial t^{2}}-2\,\frac{i}{\hbar}\,\frac{H}{R}\frac{\partial R}{\partial t}-\frac{H^{2}}{\hbar{}^{2}}-\frac{i}{\hbar}\frac{\partial H}{\partial t}$ (10)
@@
$H^{2}=m^{2}\gamma^{2}c^{4}+\hbar{}^{2}(\frac{1}{R}\frac{\partial^{2}R}{\partial t^{2}}-c^{2}\frac{\nabla{}^{2}R}{R})$ (11)
$c^{2}\frac{\vec{\nabla}R}{R}\cdot\vec{p}+\frac{H}{R}\,\frac{\partial R}{\partial t}+\frac{1}{2}\frac{\partial H}{\partial t}=0$ (12)
Now we assume that eq. (3), with $m\gamma$ instead of m, still holds (we will justify later this continuity equation). By employing $\vec{\nabla}\cdot\vec{v}=0$, we get just the expansion of $\frac{dR^{2}}{dt}$. Of course, we can refer it simply to R:
$\frac{\partial R}{\partial t}+\vec{v}\cdot\vec{\nabla}R=\frac{dR}{dt}=0$ (13)
as well as to any function of R. Indeed, since $\vec{v}$ is constant, eq. (13) also holds for any function of any-order derivative of $R$ over time or space, and thus for $V_{Q}$ and H. We are therefore assuming $\frac{dH}{dt}=0$, but with the possibility that $\frac{\partial H}{\partial t}\neq 0$, and the same for $V_{Q}$. Substituting eq. (13) into eq. (12) one finds:
$\frac{1}{R}\frac{\partial R}{\partial t}(H-m\gamma c^{2})+\frac{1}{2}\frac{\partial H}{\partial t}=0$ (14)
@@ III The two-time Bohmian model
In the first place it is therefore necessary to admit that, even in the case of a free particle, $\vec{p}$ is not constant. Indeed, inside the Heisenberg picture of the standard interpretation, Schrödinger and Dirac recognized that an omnidirectional velocity - which we will call $v_{i}$, with $i$ standing for intrinsic - of a sinusoidal motion of frequency $\frac{2m\gamma c{}^{2}}{\hbar}\,$ and amplitude about $\frac{\hbar}{2m\gamma c}\,$, known as Zitterbewegung, must be superimposed on the actually observed velocity in average value [34, 35]. This motion has been much commented on, sometimes related to spin [41, 42, 43, 44] and Quantum Potential [45], but the debate on its exact interpretation, and even observability, does not seem yet resolved [36, 37, 38, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. Hopefully, to rediscover this peculiar movement inside a deterministic theory could clarify its nature. It can only be due to the Quantum Potential; but equation of motion (6) still seems inadequate: in fact it has been shown that putting $\frac{dp_{i}}{dt}=-\mid\vec{\nabla}V_{Q}\mid$ leads to inconsistency in [1 + 1] dimensions [40].
@@
$v_{c_{x}}\equiv\frac{\partial X}{\partial t}\quad\quad v_{i_{x}}\equiv\frac{\partial X}{\partial\tau}\,$ (16)
having so: $dX=v_{c_{x}}dt+v_{i_{x}}d\tau$. If we suppose that $v_{c_{x}}$ and $v_{i_{x}}$ depend, respectively, only on $t$ and $\tau$, we have:
[…]
@@
[…]
@@
$m\gamma_{c}c^{2}+V=const\equiv E_{c}$ (19)
@@
$\frac{\partial R^{2}}{\partial t}+\vec{\nabla}\cdot(R^{2}{\vec{v_{c}})}=0$ (21)
@@
$\frac{\partial R}{\partial\tau}-v_{i_{s}}{\frac{dR}{ds}}=0$ (22)
@@
$V_{Q}+\frac{1}{2}m\gamma_{c}v_{i_{s}}^{2}\equiv E_{q_{s}}$ (23)
@@
$m\gamma_{c}\frac{dv_{i_{s}}}{d\tau}=-\frac{dV_{Q}}{ds}$ (24)
@@
Certainly, guidance equation $\vec{p}_{c}=\hbar\,Im\,(\frac{\vec{\nabla}\psi}{\psi})$ is still obtained by taking the gradient of $\psi$ and applying $\vec{\nabla}S=\vec{p}_{c}$.
@@ IV Zitterbewegung in the new model
$\frac{\partial F}{\partial t}+\vec{v_{c}}\cdot\vec{\nabla}F=\frac{dF}{dt}=0$ (25)
@@
$\frac{\partial F}{\partial t}+{v_{c}}\dot{F}=0$ (26)
@@
$\frac{\partial^{2}F}{\partial t^{2}}-{v_{c}^{2}}\ddot{F}=0$ (27)
@@
$H=\pm\sqrt{m^{2}\gamma_{c}^{2}c^{4}-\hbar{}^{2}\frac{c^{2}}{\gamma_{c}^{2}}\frac{\ddot{R}}{R}}$ (28)
$v_{c}\,\frac{\dot{R}}{R}\,(m\,\gamma_{c}\,c^{2}-H)+\frac{1}{2}\frac{\partial H}{\partial t}=0$ (29)
@@
$v_{c}\,\frac{\dot{R}}{R}\,(\sqrt{\beta}-\beta)=-\frac{1}{4}\frac{\partial\beta}{\partial t}$ (30)
@@
$\frac{\dot{R\,}}{R}\,=\frac{1}{4}\,\frac{\dot{\beta}}{\sqrt{\beta}-\beta}$ (31)
@@
$1+\frac{c_{1}}{R^{2}}=\sqrt{\beta}$ (32)
@@
$(1+\frac{c_{1}}{R^{2}})^{2}=1-\lambda^{2}\,\frac{\ddot{R}}{R}$ (34)
@@
$\dot{R}=\pm\frac{1}{\lambda}\,\sqrt{\frac{c_{1}^{2}}{R^{2}}-4c_{1}ln\,R+c_{2}}$ (35)
@@
$\dot{R}=\pm\frac{\sqrt{c_{1}}}{\lambda}\,\sqrt{\frac{c_{1}}{R^{2}}-\frac{c_{1}}{R_{M}^{2}}+4\,ln\,\frac{R_{M}}{R}}$ (36)
@@
$\int\frac{dR}{\sqrt{f(\frac{R_{M}^{2}}{R^{2}}-1)+4\,ln\,\frac{R_{M}}{R}}}=-\frac{R_{M}\sqrt{f}}{\lambda}\mid x-x_{M}\mid+c_{3}$ (38)
@@
$\sqrt{1-\frac{R}{R_{M}}}\simeq\frac{\sqrt{f(f+2)}}{\sqrt{2}\lambda}\,\mid x-x_{M}\mid+c_{4}$ (39)
@@
$R\simeq R_{M}(1-\frac{f(f+2)}{2\lambda^{2}}(x-x_{M})^{2})$ (40)
@@
$\simeq\frac{1}{R_{M}^{2}}(1+\frac{f(f+2)}{\lambda^{2}}(x-x_{M})^{2})$ (41)
@@
$V_{Q}\simeq m\gamma_{c}c^{2}f\,(1+\frac{f(f+2)}{\lambda^{2}}(x-x_{M})^{2})$ (42)
@@
$X(t,\tau)\simeq X_{t_{0}}+X_{\tau_{0}}+v_{c}\,t+\Delta X\,sin(\frac{c}{\lambda}\,f\sqrt{2(f+2)})\tau)$ (43)
$v_{i_{x}}(\tau)\simeq\Delta X\,\frac{c}{\lambda}\,f\sqrt{2(f+2)})\,cos(\frac{c}{\lambda}\,f\sqrt{2(f+2)})\tau)$ (44)
@@
$\frac{\lambda}{f\sqrt{2(f+2)}}\times m\gamma_{c}c=\frac{\hbar}{f\sqrt{2(f+2)}\gamma_{c}}\sim\frac{\hbar}{\gamma_{c}}$ (45)
@@
$\Delta X=\frac{c\sqrt{2}\sqrt{1-\sqrt{1-\beta^{2}}}}{\omega_{0}(1-\beta^{2})^{\frac{1}{4}}}$ (46)
where $\omega_{0}$ is the non-relativistic pulsation and $\beta=\frac{v_{i_{x_{MAX}}}}{c}$. But this is incompatible with our model: by imposing $v_{i_{x_{MAX}}}=c$, we obtain $DX\rightarrow\infty$, therefore not being able to represent the minimum positional uncertainty.
@@
$\frac{1}{2}m\gamma_{c}c^{2}\times\frac{\pi\lambda}{f\sqrt{2(f+2)\,}c}=\frac{\pi\hbar}{2\,f\sqrt{2(f+2)}\gamma_{c}}\sim\frac{\hbar}{\gamma_{c}}$ (47)
@@
$E_{q}=V_{Qm}+\frac{1}{2}m\gamma_{c}c{}^{2}=m\gamma_{c}c^{2}(f+\frac{1}{2})$ (48)
@@
It is noteworthy that Zitterbewegung in $\tau$-time is able to justify the physical origin of the phase to be associated with each path in Feynman’s formulation of Quantum Mechanics. The approach has been specified in [67], through a particle’s oscillation forward and backward in t-time. In this model, that type of motion arises spontaneously: it is the found Zitterbewegung in $\tau$, represented by the wavefunction extra phase factor $e^{\frac{i}{\hbar}S_{\tau}}$.

### v18

@@ II Relativistic generalization issue for the quantum potential in case of constant momentum
$\frac{\partial S}{\partial t}+\frac{(\vec{\nabla}S){}^{2}}{2m}\,+V-\frac{\hbar{}^{2}}{2m}\,\frac{\nabla{}^{2}R}{R}=0,$ (2)
$\frac{\partial R^{2}}{\partial t}+\vec{\nabla}\cdot(R^{2}\frac{\vec{\nabla}S}{m})=0.$ (3)
@@
$H=\frac{p{}^{2}}{2m}+V-\frac{\hbar{}^{2}}{2m}\,\frac{\nabla{}^{2}R}{R},$ (4)
@@
$V_{Q}\equiv-\frac{\hbar{}^{2}}{2m}\,\frac{\nabla{}^{2}R}{R}.$ (5)
@@
$\frac{{\rm{d}}\vec{p}}{{\rm{d}}t}=-\vec{\nabla}V-\vec{\nabla}V_{Q}.$ (6)
@@
$\frac{\nabla{}^{2}\psi}{\psi}=\frac{\nabla{}^{2}R}{R}+2\,\frac{\rm{i}}{\hbar}\,\frac{\vec{\nabla}R}{R}\,\cdot\,\vec{p}-\frac{p^{2}}{\hbar{}^{2}},$ (7)
@@
$\nabla{}^{2}\psi-\frac{1}{c^{2}}\,\frac{\partial^{2}\psi}{\partial t^{2}}=\frac{m{}^{2}c^{2}}{\hbar{}^{2}}\,\psi,$ (8)
@@
$\frac{\nabla{}^{2}R}{R}+2\,\frac{\rm{i}}{\hbar}\,\frac{\vec{\nabla}R}{R}\,\cdot\,\vec{p}=\frac{m{}^{2}\gamma^{2}c^{2}}{\hbar{}^{2}}+\frac{1}{c^{2}\psi}\,\frac{\partial^{2}\psi}{\partial t^{2}},$ (9)
@@
$\frac{1}{\psi}\frac{\partial^{2}\psi}{\partial t^{2}}=\frac{1}{R}\frac{\partial^{2}R}{\partial t^{2}}-2\,\frac{\rm{i}}{\hbar}\,\frac{H}{R}\frac{\partial R}{\partial t}-\frac{H^{2}}{\hbar{}^{2}}-\frac{\rm{i}}{\hbar}\frac{\partial H}{\partial t}.$ (10)
@@
$H^{2}=m^{2}\gamma^{2}c^{4}+\hbar{}^{2}(\frac{1}{R}\frac{\partial^{2}R}{\partial t^{2}}-c^{2}\frac{\nabla{}^{2}R}{R}),$ (11)
$c^{2}\frac{\vec{\nabla}R}{R}\cdot\vec{p}+\frac{H}{R}\,\frac{\partial R}{\partial t}+\frac{1}{2}\frac{\partial H}{\partial t}=0.$ (12)
Now we assume that eq. (3), with $m\gamma$ instead of m, still holds (we will justify later this continuity equation). By employing $\vec{\nabla}\cdot\vec{v}=0$, we get just the expansion of $\frac{{\rm{d}}R^{2}}{{\rm{d}}t}$. Of course, we can refer it simply to R:
$\frac{\partial R}{\partial t}+\vec{v}\cdot\vec{\nabla}R=\frac{{\rm{d}}R}{{\rm{d}}t}=0,$ (13)
as well as to any function of R. Indeed, since $\vec{v}$ is constant, eq. (13) also holds for any function of any-order derivative of $R$ over time or space, and thus for $V_{Q}$ and H. We are therefore assuming $\frac{{\rm{d}}H}{{\rm{d}}t}=0$, but with the possibility that $\frac{\partial H}{\partial t}\neq 0$, and the same for $V_{Q}$. Substituting eq. (13) into eq. (12) one finds:
$\frac{1}{R}\frac{\partial R}{\partial t}(H-m\gamma c^{2})+\frac{1}{2}\frac{\partial H}{\partial t}=0,$ (14)
@@ III The two-time Bohmian model
In the first place it is therefore necessary to admit that, even in the case of a free particle, $\vec{p}$ is not constant. Indeed, inside the Heisenberg picture of the standard interpretation, Schrödinger and Dirac recognized that an omnidirectional velocity - which we will call $v_{i}$, with $i$ standing for intrinsic - of a sinusoidal motion of frequency $\frac{2m\gamma c{}^{2}}{\hbar}\,$ and amplitude about $\frac{\hbar}{2m\gamma c}\,$, known as Zitterbewegung, must be superimposed on the actually observed velocity in average value [34, 35]. This motion has been much commented on, sometimes related to spin [41, 42, 43, 44] and Quantum Potential [45], but the debate on its exact interpretation, and even observability, does not seem yet resolved [36, 37, 38, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. Hopefully, to rediscover this peculiar movement inside a deterministic theory could clarify its nature. It can only be due to the Quantum Potential; but equation of motion (6) still seems inadequate: in fact it has been shown that putting $\frac{{\rm{d}}p_{i}}{{\rm{d}}t}=-\mid\vec{\nabla}V_{Q}\mid$ leads to inconsistency in [1 + 1] dimensions [40].
@@
$v_{c_{x}}\equiv\frac{\partial X}{\partial t}\quad\quad v_{i_{x}}\equiv\frac{\partial X}{\partial\tau},$ (16)
having so: ${\rm{d}}X=v_{c_{x}}{\rm{d}}t+v_{i_{x}}{\rm{d}}\tau$. If we suppose that $v_{c_{x}}$ and $v_{i_{x}}$ depend, respectively, only on $t$ and $\tau$, we have:
$X(t,\tau)=X_{t_{0}}+X_{\tau_{0}}+\intop_{0}^{t}v_{c_{x}}(\xi)\,{\rm{d}}\xi+\intop_{0}^{\tau}v_{i_{x}}(\zeta)\,{\rm{d}}\zeta,$ (17)
@@
$S_{t}=\int{(-\frac{mc^{2}}{\gamma_{c}}-V-V_{Q}){\rm{d}}t},\ S_{\tau}=\int(\frac{1}{2}m\gamma_{c}v_{i_{s}}^{2}-V_{Q}){\rm{d}}\tau$, where $v_{i_{s}}$ is the intrinsic velocity in the generic direction $\hat{s}$. Here we have to point out in advance that, on respect to two different spatial directions $\hat{s}$ and $\hat{u}$, it will result $v_{i_{s}}\neq v_{i_{u}}$. Therefore $\psi$ depends on $s$ through $S_{\tau}$ (it is actually a $\psi_{s}$) and hence the particle comes to be represented with a different wavefunction for each direction of space. Of course, we will demand that the meaning of $\psi$, and in particular of $R$, be the same as in the standard case.
@@
$m\gamma_{c}c^{2}+V=const\equiv E_{c},$ (19)
@@
$\frac{\partial R^{2}}{\partial t}+\vec{\nabla}\cdot(R^{2}{\vec{v_{c}})}=0.$ (21)
@@
$\frac{\partial R}{\partial\tau}-v_{i_{s}}{\frac{{\rm{d}}R}{{\rm{d}}s}}=0,$ (22)
@@
$V_{Q}+\frac{1}{2}m\gamma_{c}v_{i_{s}}^{2}\equiv E_{q_{s}},$ (23)
@@
$m\gamma_{c}\frac{{\rm{d}}v_{i_{s}}}{{\rm{d}}\tau}=-\frac{{\rm{d}}V_{Q}}{{\rm{d}}s},$ (24)
@@
Certainly, guidance equation $\vec{p}_{c}=\hbar\,{\rm{Im}}\,(\frac{\vec{\nabla}\psi}{\psi})$ is still obtained by taking the gradient of $\psi$ and applying $\vec{\nabla}S=\vec{p}_{c}$.
@@ IV Zitterbewegung in the new model
$\frac{\partial F}{\partial t}+\vec{v_{c}}\cdot\vec{\nabla}F=\frac{{\rm{d}}F}{{\rm{d}}t}=0,$ (25)
@@
$\frac{\partial F}{\partial t}+{v_{c}}\dot{F}=0,$ (26)
@@
$\frac{\partial^{2}F}{\partial t^{2}}-{v_{c}^{2}}\ddot{F}=0.$ (27)
@@
$H=\pm\sqrt{m^{2}\gamma_{c}^{2}c^{4}-\hbar{}^{2}\frac{c^{2}}{\gamma_{c}^{2}}\frac{\ddot{R}}{R}},$ (28)
$v_{c}\,\frac{\dot{R}}{R}\,(m\,\gamma_{c}\,c^{2}-H)+\frac{1}{2}\frac{\partial H}{\partial t}=0.$ (29)
@@
$v_{c}\,\frac{\dot{R}}{R}\,(\sqrt{\beta}-\beta)=-\frac{1}{4}\frac{\partial\beta}{\partial t},$ (30)
@@
$\frac{\dot{R\,}}{R}\,=\frac{1}{4}\,\frac{\dot{\beta}}{\sqrt{\beta}-\beta}.$ (31)
@@
$1+\frac{c_{1}}{R^{2}}=\sqrt{\beta},$ (32)
@@
$(1+\frac{c_{1}}{R^{2}})^{2}=1-\lambda^{2}\,\frac{\ddot{R}}{R}.$ (34)
@@
$\dot{R}=\pm\frac{1}{\lambda}\,\sqrt{\frac{c_{1}^{2}}{R^{2}}-4c_{1}{\rm{ln}}\,R+c_{2}},$ (35)
@@
$\dot{R}=\pm\frac{\sqrt{c_{1}}}{\lambda}\,\sqrt{\frac{c_{1}}{R^{2}}-\frac{c_{1}}{R_{M}^{2}}+4\,{\rm{ln}}\,\frac{R_{M}}{R}}.$ (36)
@@
$\int\frac{{\rm{d}}R}{\sqrt{f(\frac{R_{M}^{2}}{R^{2}}-1)+4\,{\rm{ln}}\,\frac{R_{M}}{R}}}=-\frac{R_{M}\sqrt{f}}{\lambda}\mid x-x_{M}\mid+c_{3},$ (38)
@@
$\sqrt{1-\frac{R}{R_{M}}}\simeq\frac{\sqrt{f(f+2)}}{\sqrt{2}\lambda}\,\mid x-x_{M}\mid+c_{4},$ (39)
@@
$R\simeq R_{M}(1-\frac{f(f+2)}{2\lambda^{2}}(x-x_{M})^{2}),$ (40)
@@
$\simeq\frac{1}{R_{M}^{2}}(1+\frac{f(f+2)}{\lambda^{2}}(x-x_{M})^{2}).$ (41)
@@
$V_{Q}\simeq m\gamma_{c}c^{2}f\,(1+\frac{f(f+2)}{\lambda^{2}}(x-x_{M})^{2}).$ (42)
@@
$X(t,\tau)\simeq X_{t_{0}}+X_{\tau_{0}}+v_{c}\,t+\Delta X\,{\rm{sin}}(\frac{c}{\lambda}\,f\sqrt{2(f+2)})\tau),$ (43)
$v_{i_{x}}(\tau)\simeq\Delta X\,\frac{c}{\lambda}\,f\sqrt{2(f+2)})\,{\rm{cos}}(\frac{c}{\lambda}\,f\sqrt{2(f+2)})\tau),$ (44)
@@
$\frac{\lambda}{f\sqrt{2(f+2)}}\times m\gamma_{c}c=\frac{\hbar}{f\sqrt{2(f+2)}\gamma_{c}}\sim\frac{\hbar}{\gamma_{c}},$ (45)
@@
$\Delta X=\frac{c\sqrt{2}\sqrt{1-\sqrt{1-\beta^{2}}}}{\omega_{0}(1-\beta^{2})^{\frac{1}{4}}},$ (46)
where $\omega_{0}$ is the non-relativistic pulsation and $\beta=\frac{v_{i_{x_{MAX}}}}{c}$. But this is incompatible with our model: by imposing $v_{i_{x_{MAX}}}=c$, we obtain $\Delta X\rightarrow\infty$, therefore not being able to represent the minimum positional uncertainty.
@@
$\frac{1}{2}m\gamma_{c}c^{2}\times\frac{\pi\lambda}{f\sqrt{2(f+2)\,}c}=\frac{\pi\hbar}{2\,f\sqrt{2(f+2)}\gamma_{c}}\sim\frac{\hbar}{\gamma_{c}}.$ (47)
@@
$E_{q}=V_{Qm}+\frac{1}{2}m\gamma_{c}c{}^{2}=m\gamma_{c}c^{2}(f+\frac{1}{2}),$ (48)
@@
It is noteworthy that Zitterbewegung in $\tau$-time is able to justify the physical origin of the phase to be associated with each path in Feynman’s formulation of Quantum Mechanics. The approach has been specified in [67], through a particle’s oscillation forward and backward in t-time. In this model, that type of motion arises spontaneously: it is the found Zitterbewegung in $\tau$, represented by the wavefunction extra phase factor ${\rm{e}}^{\frac{{\rm{i}}}{\hbar}S_{\tau}}$.